# AgentMove: Predicting Human Mobility Anywhere Using Large Language Model based Agentic Framework

### v1

@@ Abstract
Human mobility prediction plays a crucial role in various real-world applications. Although deep learning based models have shown promising results over the past decade, their reliance on extensive private mobility data for training and their inability to perform zero-shot predictions, have hindered further advancements. Recently, attempts have been made to apply large language models (LLMs) to mobility prediction task. However, their performance has been constrained by the absence of a systematic design of workflow. They directly generate the final output using LLMs, which limits the potential of LLMs to uncover complex mobility patterns and underestimates their extensive reserve of global geospatial knowledge. In this paper, we introduce AgentMove, a systematic agentic prediction framework to achieve generalized mobility prediction for any cities worldwide. In AgentMove, we first decompose the mobility prediction task into three sub-tasks and then design corresponding modules to complete these subtasks, including spatial-temporal memory for individual mobility pattern mining, world knowledge generator for modeling the effects of urban structure and collective knowledge extractor for capturing the shared patterns among population. Finally, we combine the results of three modules and conduct a reasoning step to generate the final predictions. Extensive experiments on mobility data from two sources in 12 cities demonstrate that AgentMove outperforms the best baseline more than 8% in various metrics and it shows robust predictions with various LLMs as base and also less geographical bias across cities. Codes and data can be found in https://github.com/tsinghua-fib-lab/AgentMove.
Introduction
@@
Mobility prediction is of great importance in many real-world scenarios, e.g., recommend travel services, pre-activating mobile applications for potential usage, seamless switching of cellular network signals and efficient traffic management. In the recent years, deep learning based models (Liu et al. 2016; Wu et al. 2017; Feng et al. 2018; Yang et al. 2020; Yang, Liu, and Zhao 2022) are widely applied and has achieved good results due to its advantages in capturing the high-order transition capabilities and mining the shared mobility patterns between users. However, existing works have several key drawbacks. First, the success of deep learning models rely on the collection of large amounts of private mobility data. Second, the trained model is difficult to be applied in the zero-shot mobility prediction settings. Finally, the prediction accuracy is still lower due to the limited sequential modelling capability of small deep learning models and lack of deeply understanding of common sense in human daily life and urban structures.
@@
Recently, large language models (LLMs) have made significant progress, achieving advanced results far surpassing previous methods in areas such as dialogue based role-playing, code generation and testing, and mathematical problem solving. In the field of spatial temporal data mining, researchers also explore the potential of applying LLMs many real-world tasks, such as time series forecasting (Gruver et al. 2024; Li et al. 2024b), travel planning (Xie et al. 2024; Li et al. 2024a), trajectory analysis (Luo et al. 2024; Zhang et al. 2023), and so on. Furthermore, several recent works (Wang et al. 2023; Beneduce, Lepri, and Luca 2024) investigate the feasibility of employing the LLMs as the base model of mobility predictor for addressing the previous limitations of deep learning based models and achieve promising results. These works first convert the trajectory to a language based sentence and then utilize the powerful sequential modelling capacities of LLMs to directly output the mobility prediction results. However, due to the absence of a systematic design of whole procedure, they ignore the crucial components of human mobility modeling and thus achieve limited performance. In summary, these methods fail to effectively capture the complicated individual mobility pattern, ignore modeling the effects of urban structure and discovering the shared mobility patterns among population.
@@
In this paper, we propose AgentMove, a systematic agentic framework for generalized mobility prediction in any cities around the world. In AgentMove, by integrating the domain knowledge of human mobility, we implement the core components in the general agentic framework (Wang et al. 2024; Xi et al. 2023), including the planning module, memory module, world knowledge module, external tool module and reasoning module. For the planning module in AgentMove, we introduce a manually designed mobility prediction task decomposition module by considering the most important factors for mobility prediction. After decomposition, we generate three sub-tasks: individual mobility pattern mining, shared mobility pattern discovering and urban structure modelling. Firstly, we implement a spatial-temporal memory module for individual mobility pattern mining. It contains three submodules, short-term memory, long-term memory and user profiles, to capture the multi-level mobility patterns of individual. Compared with the pure LLM methods, the memory module makes AgentMove to memorize the past mobility history and continuously learn from the experiences efficiently. Secondly, we design a world knowledge generator to explicitly extract the inherent geospatial knowledge of the urban structure from LLMs to help modelling the effects of multi-scale urban structure on the human mobility especially the exploration behavior of human mobility. Thirdly, we endow AgentMove with the ability of discovering the shared mobility patterns from various trajectories of different users by a collective knowledge extractor. It utilizes the NetworkX as a external tool to organize the trajectories from various users to form a location transition graph and then extract important neighbor locations for prediction. Finally, we combine all the results from all the modules modules and perform a final reasoning step to generate the prediction. In summary, our contributions can be summarized as follows,
@@
  • To our best knowledge, this is the first attempt to apply LLM-based agentic framework into the field of mobility prediction. We build an effective mobility prediction framework by considering the crucial characteristics of human mobility into the designs of core components.
@@
  • We design a mobility prediction task decomposition module and integrate the modelling of different aspects of human mobility into different modules in AgentMove. Specifically, it contains a spatial-temporal memory module for individual mobility pattern mining, a world knowledge generator for modeling effects of urban structures and a graph based collective knowledge extractor for discovering the shared mobility patterns among population.
@@
  • Extensive experiments on mobility trajectories from 2 sources in 12 cities demonstrate the effectiveness of proposed AgentMove, which outperforms the best baseline with a performance increase of over 8% in the majority of cases. Additionally, AgentMove presents superior adaptability to different LLMs, as well as greater stability and reduced bias in prediction results across various cities worldwide.
Preliminaries
@@
Here, we define the mobility prediction task and related concepts, which are used in the following section.
Figure 1: Different paradigms of deep learning model based mobility predictor and LLM based mobility predictor
Definition 1 (POI)
@@
A POI point $p\in P$ is represented as a tuple $\langle id,cate,lon,lat,addr\rangle$, where $id$ is the unique identifier, $cate$ is the category (e.g., restaurant), $lon$ and $lat$ are the coordinates of the POI, $addr$ is the text address of POI.
@@
A trajectory of user $u\in U$ is represented as $T_{u}=\{(p_{1},t_{1}),(p_{2},t_{2}),\dots,(p_{n},t_{n})\}$, where $p_{i}\in P$ is the $i$-th POI visited by the user and $t_{i}$ is the timestamp of the visit.
@@
Given the historical movement data $\mathcal{C}_{u}$, $\mathcal{H}_{u}$ as well as available external knowledge $\mathcal{K}$ (e.g., worldwide geospatial information), the objective is to predict the next POI $p_{n+1}$ that user $u$ will visit. Formally, this paper aims to learn a mapping function $f$:
@@
Figure 1 illustrates the differences between the deep learning based paradigm and the LLM based paradigm in the mobility prediction task. The deep learning model needs collecting training data before conduct the prediction task, which means it cannot directly used in the zero-shot scenario. LLM based method can directly applied into any scenario after carefully ‘format converter’ (known as prompt engineering).
Methods
Figure 2: The framework of AgentMove, including spatial temporal memory unit for multi-scale mobility pattern of single user, spatial world model for multi-level saptial structure of nearby locations and social world model for extracting shared mobility patterns of user group.
Overview
As shown in Figure 2, AgentMove consists of five core components: task decomposition module, spatial-temporal memory module, world knowledge generator, collective knowledge extractor and the final reasoning module. As the high-level planning module, the task decomposition module is designed to break down the whole mobility prediction task into subtasks—personalized mobility pattern mining, collective mobility pattern discovery and the effects of urban structures—by considering the crucial factors of mobility. Detailed design of other components are introduced as follows.
Spatial-temporal Memory for Personalized Multi-scale Periodicity Behavior Modelling
The spatial-temporal memory module is designed to effectively capture, store and leverage mobility patterns, providing crucial insights for the personalized and multi-scale periodicity behavior modelling in mobility prediction. Inspired by the memory design principles in general LLM-based agents (Zhang et al. 2024), our spatial-temporal memory functions through three essential processes: memory organization, memory writing, and memory reading. The whole framework of spatial–temporal memory module is presented in Figure 3.
@@ Methods
Memory Organization
The spatial-temporal memory is structured into three components to capture multifaceted nature of user mobility patterns:
  • User Profile Unit. This unit provides a summary description of the user’s mobility behavior as the user mobile profile, which offers deeper insights into when and why the user visits certain locations. The user profile is dynamically generated based on the current long-term memory introduced in the following part, allowing AgentMove to adapt to the evolving user preferences.
  • Long-term Memory Unit. This unit retains users’ long-term mobility patterns, capturing overarching trends and recurring sequences in their movement history. It functions similarly to how LLMs store long-term dependencies in textual data.
  • Short-term Memory Unit. This unit focuses on users’ recent mobility patterns, providing dynamic updates that reflect the latest movements and short-term variations.
@@
Memory Reading
The memory reading process involves generating spatial-temporal context relevant prompts from the structured memory to to enhance AgentMove’s predictive capabilities. This process consists of three key steps:
@@
World Knowledge Generator for Enhancing Exploration Behavior Modelling
Lots of research (Jiang et al. 2016) shows that individual movement typically includes two types of behaviors: return and explore. As introduced before, the return behavior has been well-captured by the spatial-temporal memory module. In this section, we introduce the world knowledge generator to extract geospatial knowledge from LLMs and construct multi-scale urban structure to enable the modelling of explore behavior of mobility. To extract geospatial knowledge effectively, we propose to alignment the knowledge of LLMs and the urban space of trajectory via text address. Once the space is alignment, we explicitly motivate the LLMs to generate the potential explore candidate places from the multi-scale urban structure view.
Alignment via Address
Many existing works (Feng et al. 2018; Luo, Liu, and Liu 2021; Lin et al. 2021; Cui et al. 2021; Qin et al. 2022; Hong et al. 2023) on mobility prediction usually represent the locations directly using latitude and longitude coordinates or discrete spatial area IDs. While this approach facilitates the easy construction of deep learning-based spatial encoding, it is not suitable for LLMs. Since LLMs are trained on large scale human-generated text, they, like human, are not inherently adept at understanding the precise coordinates (Gurnee and Tegmark 2023) or discrete area IDs. Thus, we propose to utilize the text address which human is familiar to describe the coarse location of trajectory. While text address is not precise as the coordinates, it is more natural and easy to be aligned with the existing spatial knowledge in the LLMs.
@@
Multi-scale Urban Structure
@@
Collective Knowledge Extractor for Shared Mobility Pattern Modelling
In the previous two sections, we introduce the spatial-temporal memory module and world knowledge generator for the individual level mobility modelling. Here, we introduce the collective knowledge extractor to enable capture the sharing mobility pattern among users to further improve the mobility prediction. In one world, we first construct a global location transition graph using NetworkX  [^2] https://networkx.org/documentation/stable/index.html by aggregating the location transitions from various users. We then employ a LLM to perform simple reasoning on the graph, utilizing the function in NetworkX as tool to generate potential locations visited by other users with similar mobility pattern.
Building Location Transition Graph
@@
Finally, we design prompts to employ LLM to analyze and summarize the information from different views and perform a final reasoning step to generate the prediction with reasons. The prompts for output format requirements are also be placed here to ensure that the output format meets the requirements as much as possible.
Evaluation
@@
vs Deep Learning 5.71% 45.31% 31.87% 21.05% 13.70% 16.01% -3.03% 8.96% 2.71% 8.57% 26.76% 31.08%
vs Best Baselines 5.71% 13.41% 10.70% 21.05% 7.79% 16.01% -3.03% 7.35% 2.71% 8.57% 5.88% 10.40%
@@
Table 1: The main results of baselines and AgentMove. GPT4omini is used as the base LLM for all the LLM-based methods in the table. FPMC and deep learning based models are first trained on the training set of each city and LLM-based models are directly evaluated on the test set with the zero-shot prediction settings.
Settings
Datasets
@@
We use the global Foursquare checkin data (Yang, Zhang, and Qu 2016) and recent public released ISP trajectory data (Feng et al. 2019) to conduct the experiments. The Foursquare data contains checkins from 415 cities which covers about 18 months from April 2012 to September 2013. The ISP trajectory data is from Shanghai with 325215 records, covering April 19 to April 26 in 2016. Compared with Foursquare data, ISP data is much denser and was open-sourced only two months ago  [^3] https://github.com/vonfeng/DPLink/tree/master/data, which is beyond the training period of all the LLMs used in the experiment. This ensures that the evaluation results are not affected by potential data leakage.
@@
To evaluate the general mobility prediction ability of AgentMove, we select 12 cities from the Foursquare dataset and the entire ISP trajectory data to conduct the experiments. We follow the preprocessing procedure (Hong et al. 2023; Feng et al. 2019) to process the trajectories data. Detailed description about preprocessing can refer to the appendix. We select the Tokyo, Moscow and SaoPaulo with the largest amount of Foursquare check-in data and the ISP data from Shanghai to conduct the main analysis in the experiments and results of 12 cities are discussed in the final section of experiment. Due to the cost of the various API calling, e.g., Llama3.1-405B, we randomly sample 200 instances from the testing set for each city to calculate the performance in the experiments.
Baselines
We have three kinds of baselines, Markov methods, deep learning models and LLM-based methods.
  • FPMC (Rendle, Freudenthaler, and Schmidt-Thieme 2010) It combines the matrix factorization and Markov chains methods together for sequential modelling.
  • RNN (Feng et al. 2018) It is a simple RNN based mobility prediction model as regarding the mobility sequence as general sequence.
  • DeepMove (Feng et al. 2018) It contains a LSTM for capturing the short-term sequential transition and an attention unit for extracting long-term periodical patterns.
  • LSTPM (Sun et al. 2020) It consists of a non-local network for long-term preference modeling and a geo-dilated RNN for short-term preference learning.
  • LLM-Mob (Wang et al. 2023) It is the first work to apply LLM (GPT-3.5) to predict the next location with carefully prompt engineering.
  • LLM-ZS (Beneduce, Lepri, and Luca 2024) Following LLM-Mob, it simplifies the prompts and testifies more LLMs in zero-shot mobility prediction task.
We use widely used Accuracy@1, Accuracy@5, and NDCG@5 as the main evaluation metrics (Sun et al. 2020; Luca et al. 2021) in the experiments.
Implementation

### v2

# AgentMove: A Large Language Model based Agentic Framework for Zero-shot Next Location Prediction

@@ Abstract
Next location prediction plays a crucial role in various real-world applications. Recently, due to the limitation of existing deep learning methods, attempts have been made to apply large language models (LLMs) to zero-shot next location prediction task. However, they directly generate the final output using LLMs without systematic design, which limits the potential of LLMs to uncover complex mobility patterns and underestimates their extensive reserve of global geospatial knowledge. In this paper, we introduce AgentMove, a systematic agentic prediction framework to achieve generalized next location prediction. In AgentMove, we first decompose the mobility prediction task and design specific modules to complete them, including spatial-temporal memory for individual mobility pattern mining, world knowledge generator for modeling the effects of urban structure and collective knowledge extractor for capturing the shared patterns among population. Finally, we combine the results of three modules and conduct a reasoning step to generate the final predictions. Extensive experiments utilizing mobility data from two distinct sources reveal that AgentMove surpasses the leading baseline by 3.33% to 8.57% across 8 out of 12 metrics and it shows robust predictions with various LLMs as base and also less geographical bias across cities. Our codes are available via https://github.com/tsinghua-fib-lab/AgentMove.
AgentMove: A Large Language Model based Agentic Framework for Zero-shot Next Location Prediction
Jie Feng, Yuwei Du, Jie Zhao, Yong Li Department of Electronic Engineering, BNRist, Tsinghua University, Beijing, China, {fengjie, liyong07}@tsinghua.edu.cn
1 Introduction
@@
Mobility prediction is of great importance in many real-world scenarios, e.g., recommending travel services, pre-activating mobile applications for potential usage, seamless switching of cellular network signals and efficient traffic management. Next location prediction is one of the most important task in human mobility prediction. In recent years, deep learning based models Liu et al. (2016); Wu et al. (2017); Feng et al. (2018); Yang et al. (2020, 2022) have been widely applied and have achieved promising results due to their ability to capture the high-order transition dynamics and mining shared mobility patterns among users. However, existing approaches have several key drawbacks. First, the success of deep learning models rely on the collection of large amounts of private mobility data. Second, the trained model are challenging to apply in zero-shot mobility prediction settings. Finally, the prediction accuracy remains limited due to the constrained sequential modelling capability of smaller deep learning models and a lack of deep understanding of commonsense in human daily life and urban structures.
Figure 1: The framework of AgentMove, including three key components: spatial temporal memory unit for capturing individual mobility pattern, world knowledge generator for multi-level urban structure, and collective knowledge extractor for extracting shared mobility patterns among users.
@@
Recently, large language models (LLMs) have made significant progress, achieving advanced results that far surpass previous methods in areas such as dialogue-based role-playing, code generation and testing, and mathematics problem solving. In the field of spatial-temporal data mining, researchers are exploring the potential of applying LLMs to various real-world tasks, including time series forecasting Gruver et al. (2024); Li et al. (2024b), travel planning Xie et al. (2024); Li et al. (2024a), trajectory analysis Luo et al. (2024); Zhang et al. (2023); Du et al. (2024). Furthermore, several recent works Wang et al. (2023); Beneduce et al. (2024) investigate the feasibility of using LLMs as the base model of mobility prediction, addressing the limitations of deep learning based models and achieving promising results. These works typically convert trajectories to a language based sentence and leverage the powerful sequential modelling capacities of LLMs to directly generate the mobility predictions. However, due to the lack of a systematic design throughout the entire process, they overlook the crucial components of human mobility modeling, resulting in limited performance. In summary, these methods fail to effectively capture the complex individual mobility patterns, neglect to model the effects of urban structure and do not discover the shared mobility patterns among populations.
@@
In this paper, we propose AgentMove, a systematic agentic framework for generalized mobility prediction. By integrating domain knowledge of human mobility, we implement the core components in the general agentic framework Wang et al. (2024); Xi et al. (2023), including the planning module, memory module, world knowledge module, external tool module and reasoning module. For the planning module in AgentMove, we introduce a manually designed mobility prediction task decomposition module that considers the most important factors influencing mobility prediction. This decomposition generates three sub-tasks: individual mobility pattern mining, shared mobility pattern discovery and urban structure modelling. First, we implement a spatial-temporal memory module for individual mobility pattern mining. This module contains three submodules–short-term memory, long-term memory and user profiles–to capture the multi-level mobility patterns of individuals. Compared to pure LLM methods, the memory module enables AgentMove to retain past mobility history and efficiently learn from experiences. Second, we design a world knowledge generator to explicitly extract inherent geospatial knowledge from LLMs, aiding in the modelling the effects of multi-scale urban structures on the human mobility, particularly in relation to the exploration behavior of human mobility. Third, we equip AgentMove with the capability to discover the shared mobility patterns from various user trajectories through a collective knowledge extractor. This extractor utilizes NetworkX as an external tool to organize trajectories into a global location transition graph and then extract important neighboring locations for prediction. Finally, we combine the results from all the modules and perform a final reasoning step to generate the predictions. In summary, our contributions are as follows,
@@
  • To the best of our knowledge, this is the first attempt to apply LLM-based agentic framework to the field of mobility prediction. We build an effective mobility prediction framework by incorporating the crucial characteristics of human mobility into the design of core components.
@@
  • In AgentMove, we design a spatial-temporal memory module for individual mobility pattern mining, a world knowledge generator for modeling effects of urban structures, and a graph based collective knowledge extractor for discovering the shared mobility patterns among populations.
@@
  • Extensive experiments on mobility trajectories from two sources in 12 cities demonstrate the effectiveness of proposed AgentMove, which outperforms the best baseline, achieving performance improvements ranging from 3.33% to 8.57% in most cases. Additionally, AgentMove presents superior adaptability to different LLMs, as well as greater stability and reduced bias in prediction results across various cities worldwide.
2 Preliminaries
@@
We define the mobility prediction task and related concepts for use in the following section.
Figure 2: Deep learning and LLM-based mobility predictors work in different ways. Deep learning models need to learn from training data for specific regions, while LLMs predict directly using zero-shot reasoning with its world knowledge.
Definition 1 (Location)
@@
A location point $p\in P$ is represented as a tuple $\langle id,cate,lon,lat,addr\rangle$, where $id$ is the unique identifier, $cate$ is the category (e.g., restaurant), $lon$ and $lat$ are the coordinates of the location, $addr$ is the text address of location.
@@
A trajectory of user $u\in U$ is represented as $T_{u}=\{(p_{1},t_{1}),(p_{2},t_{2}),\dots,(p_{n},t_{n})\}$, where $p_{i}\in P$ is the $i$-th location visited by the user and $t_{i}$ is the timestamp of the visit.
@@
Given the historical movement data $\mathcal{C}_{u}$, $\mathcal{H}_{u}$ as well as available external knowledge $\mathcal{K}$ (e.g., worldwide geospatial information), the objective is to predict the next location $p_{n+1}$ that user $u$ will visit. Formally, this paper aims to learn a mapping function $f$:
@@
Figure 2 illustrates the differences between the deep learning based paradigm and the LLM based paradigm in the mobility prediction task. The deep learning model needs collecting training data before conduct the prediction task, which means it cannot directly used in the zero-shot scenario. LLM based method can directly applied into any scenario after carefully ‘format converter’ (known as prompt engineering). While LLM based methods can be adapted easily to new scenarios, their effectiveness may not improve as the scenarios accumulates more data. In this way, the deep learning models with more data can achieve better performance when LLM based methods fail to improving. In this paper, we propose the LLM based agent solution AgentMove for mobility prediction task which enables the continue learning and improving of LLM based mobility predictor.
3 Methods
3.1 Overview
As shown in Figure 1, AgentMove comprises five core components: task decomposition module, spatial-temporal memory module, world knowledge generator, collective knowledge extractor and the final reasoning module. Serving as the high-level planning module, the task decomposition module is designed to break down the overall mobility prediction task into subtasks—personalized mobility pattern mining, collective mobility pattern discovery and modelling the effects of urban structures—by considering the crucial factors influencing mobility. The detailed design of the other components is introduced as follows.
3.2 Spatial-temporal Memory
The spatial-temporal memory module is designed to effectively capture, store and leverage mobility patterns, providing crucial insights for the personalized and multi-scale periodicity behavior modelling in mobility prediction. Inspired by the memory design principles in general LLM-based agents Zhang et al. (2024), our spatial-temporal memory functions through three essential processes: memory organization, memory writing, and memory reading. The whole framework of spatial–temporal memory module is presented in Figure 3.
@@ 3.2 Spatial-temporal Memory
3.2.1 Memory Organization
The spatial-temporal memory is structured into three components to capture multifaceted nature of user mobility patterns: User Profile Unit. This unit provides a summary description of the user’s mobility behavior as the user mobile profile, which offers deeper insights into when and why the user visits certain locations. The user profile is dynamically generated based on the current long-term memory introduced in the following part, allowing AgentMove to adapt to the evolving user preferences; Long-term Memory Unit. This unit retains users’ long-term mobility patterns, capturing overarching trends and recurring sequences in their movement history. It functions similarly to how LLMs store long-term dependencies in textual data; Short-term Memory Unit. This unit focuses on users’ recent mobility patterns, providing dynamic updates that reflect the latest movements and short-term variations.
@@ 3.2.2 Memory Writing
3.2.3 Memory Reading
The memory reading process involves generating spatial-temporal context relevant prompts from the structured memory to enhance AgentMove’s predictive capabilities. This process consists of three key steps:
@@ 3.2.3 Memory Reading
3.3 World Knowledge Generator
Numerous studies Jiang et al. (2016) indicate that individual movement typically encompasses two types of behaviors: returning and exploring. As introduced before, the return behavior has been well-captured by the spatial-temporal memory module. In this section, we introduce the world knowledge generator, which extracts geospatial knowledge from LLMs and constructs a multi-scale urban structure to enable the modelling of explore behavior in mobility. To extract geospatial knowledge effectively, we propose aligning the knowledge of LLMs and the urban space of trajectory via text addresses. Once the spaces are aligned, we explicitly prompt the LLMs to generate potential candidate places for exploration from the perspective of the multi-scale urban structure.
3.3.1 Alignment via Address
Many existing works Feng et al. (2018); Luo et al. (2021); Lin et al. (2021); Cui et al. (2021); Qin et al. (2022); Hong et al. (2023) on mobility prediction usually represent the locations directly using latitude and longitude coordinates or discrete spatial area IDs. While this approach facilitates the easy construction of deep learning-based spatial encoding, it is not suitable for LLMs. Since LLMs are trained on large scale human-generated text, they, like human, are not inherently adept at understanding the precise coordinates Gurnee and Tegmark (2023) or discrete area IDs. Thus, we propose to utilize the text address which human is familiar to describe the coarse location of trajectory. While text address is not precise as the coordinates, it is more natural and easy to be aligned with the existing spatial knowledge in the LLMs.
@@ 3.3.1 Alignment via Address
3.3.2 Multi-scale Urban Structure
@@
3.4 Collective Knowledge Extractor
In the previous two sections, we introduce the spatial-temporal memory module and world knowledge generator for the individual-level mobility modelling. Here, we introduce the collective knowledge extractor, which captures shared mobility patterns among users to further enhance the mobility predictions. First, we construct a global location transition graph using NetworkX  [^2] https://networkx.org/documentation/stable/index.html by aggregating the location transitions from various users. We then employ a LLM to perform simple reasoning on the graph, utilizing functions in NetworkX as tool to generate potential locations visited by other users with similar mobility patterns.
3.4.1 Building Location Transition Graph
@@ 3.4.2 Reasoning on Graph
3.5 Summarization and Prediction
Finally, we design prompts to employ LLM to analyze and summarize the information from different views and perform a final reasoning step to generate the prediction with reasons. The prompts for output format requirements are also be placed here to ensure that the output format meets the requirements as much as possible. Detailed prompts can refer to the appendix.
4 Evaluation
@@
vs Deep Learning -9.76% 3.33% 4.42% 25.71% 4.27% 9.03% -11.11% -3.95% -1.49% 8.57% 40.63% 31.08%
vs Best Baselines -9.76% 3.33% 4.42% 4.55% 4.27% 0.31% -11.11% -3.95% -1.49% 8.57% 5.88% 6.82%
@@
Table 1: The main results of baselines and AgentMove. GPT4omini is used as the base LLM for all the LLM-based methods in the table. Deep learning methods are first trained on the training set of each city and LLM-based models are directly evaluated on the test set with the zero-shot prediction settings.
4.1 Settings
4.1.1 Datasets
@@
We use the global Foursquare checkin data Yang et al. (2016) and recent public released ISP GPS trajectory data Feng et al. (2019) to conduct the experiments. The Foursquare data contains checkins from 415 cities which covers about 18 months from April 2012 to September 2013. The ISP GPS trajectory data is from the mobile network logs in Shanghai with 325215 records, covering April 19 to April 26 in 2016. Compared with Foursquare data, ISP data is much denser and was open-sourced in June 2024  [^3] https://github.com/vonfeng/DPLink/tree/master/data, which is beyond the training period of all the LLMs used in the experiment. This ensures that the evaluation results are not affected by potential data leakage.
To evaluate the general mobility prediction ability of AgentMove, we select 12 cities from the Foursquare dataset and the entire ISP trajectory data to conduct the experiments. We follow the preprocessing procedure Hong et al. (2023); Feng et al. (2019) to process the trajectories data. For Foursquare checkin data, we divide each trajectory dataset into training, validation, and test sets in a ratio of 7:1:2. While the ISP data lasts only 7 days, we split the whole data into training set, validation set and testing data in a ratio of 4:1:5 for preserving enough testing data. Detailed description about preprocessing can refer to the appendix. We follow the data license in the original paper and use these trajectory data only for academic purpose.
@@
We select Tokyo, Moscow and SaoPaulo with the largest amount of Foursquare check-in data and the ISP data from Shanghai to conduct the main analysis in the experiments and results of 12 cities are discussed in the final section of experiment. We divide each trajectory dataset into training, validation, and test sets. The training and validation sets are only used to train the deep learning model, and the resulting models are compared with the LLM-based methods on the test set. Due to the cost of the various API calling, e.g., Llama3.1-405B, we randomly sample 200 instances from the testing set for each city to calculate the performance in the experiments.
4.1.2 Baselines
We compare proposed models with following baselines: FPMC Rendle et al. (2010), five deep learning models (RNN Feng et al. (2018), DeepMove Feng et al. (2018), LSTPM Sun et al. (2020), GETNext Yang et al. (2022), STHGCN Yan et al. (2023)) and three LLM-based methods(LLM-Mob Wang et al. (2023), LLM-ZS Beneduce et al. (2024), LLM-Move Feng et al. (2024c)). We use widely used Accuracy@1, Accuracy@5, and NDCG@5 as the main evaluation metrics Sun et al. (2020); Luca et al. (2021) in the experiments.
4.1.3 Implementation